OpenStudy (quickstudent):

Write an expression to represent the total volume of two different sized boxes as a sum of cubes if one of the boxes has sides with a length of 1 foot and the other has sides with a length of x feet.

OpenStudy (quickstudent):

This is not from my homework, I was just wanting an explanation of how to do this.

OpenStudy (anonymous):

\[1+x^3\] is my guess

OpenStudy (quickstudent):

Hmmm... any other guesses?

OpenStudy (quickstudent):

@iambatman ?

OpenStudy (quickstudent):

@Whitemonsterbunny17 ?

OpenStudy (quickstudent):

@zepdrix ?

OpenStudy (quickstudent):

@zaibali.qasmi ?

OpenStudy (quickstudent):

@dan815 ?

OpenStudy (quickstudent):

@study100 ?

OpenStudy (quickstudent):

@KlOwNlOvE ?

OpenStudy (quickstudent):

@midhun.madhu1987 ?

OpenStudy (quickstudent):

@HELP!!!! ?

OpenStudy (quickstudent):

@agreene ?

OpenStudy (quickstudent):

@juanpabloJR ?

OpenStudy (quickstudent):

@Mandre ?

OpenStudy (quickstudent):

@jenniferjuice ?

zepdrix (zepdrix):

We're assuming that the box has square faces? Meaning that the box is a cube? Meaning that all the sides are the same length? Then yes, I would agree with satellite. Volume of the first box, \(\Large\rm v_1=1\cdot1\cdot1=1\) Volume of the second box, \(\Large\rm v_2=x\cdot x\cdot x=x^3\) Total volume being: \(\Large\rm v_1+v_2=1+x^3\)